# Independent sets of non-geometric lattices and the maximal adjoint

### v2

@@ 6.4. Theorem
$1=rank_{\mathcal{M}^{*}}(C_{1}^{c}\cap C_{2}^{c}\cap C_{3}^{c})=rank_{\mathcal{M}^{*}}(C_{1}^{c}\cap C_{2}^{c})=rank_{\mathcal{M}^{*}}(C_{1}^{c}\cap C_{3}^{c})=rank_{\mathcal{M}^{*}}(C_{2}^{c}\cap C_{3}^{c})$
A set of size three in $\epsilon(2^{C(\mathcal{M})}\setminus\mathcal{S}(\mathcal{M}))$ is of the form $\left\{C_{2},C_{3},C_{4}\right\}$ with some $C_{1}\in\mathcal{C}(\mathcal{M})$ such that $\left\{C_{1},C_{2},C_{3}\right\},\left\{C_{1},C_{2},C_{4}\right\}\in 2^{C(\mathcal{M})}\setminus\mathcal{S}(\mathcal{M})$, which is again in $2^{C(\mathcal{M})}\setminus\mathcal{S}(\mathcal{M})$. As we also have $(\uparrow(2^{C(\mathcal{M})}\setminus\mathcal{S}(\mathcal{M})))_{3}=(2^{C(\mathcal{M})}\setminus\mathcal{S}(\mathcal{M}))_{3}$ and the maximal $rank_{\mathcal{M}^{*}}$ of a set is $3$ we get $D(\delta\mathcal{M})=2^{C(\mathcal{M})}\setminus\mathcal{S}(\mathcal{M})$,
@@ 6.5. Theorem
$n-k-rank_{\mathcal{M}^{*}}(\cap_{A\in\mathcal{A}_{i}}A)=\left|\left(\cup_{A\in\mathcal{A}_{i}}A\right)\setminus C\right|+1<\left|\mathcal{A}_{i}\right|$
@@
$n-k-rank_{\mathcal{M}^{*}}(\cap_{A\in\left(\mathcal{A}_{1}\cup\mathcal{A}_{2}\right)\setminus\left\{C\right\}}A)\leq n-k-rank_{\mathcal{M}^{*}}(\cap_{A\in\mathcal{A}_{1}\cup\mathcal{A}_{2}}A)$
$=\left|\left(\cup_{A\in\mathcal{A}_{1}\cup\mathcal{A}_{2}}A\right)\setminus C\right|+1\leq\left|\left(\cup_{A\in\mathcal{A}_{1}}A\right)\setminus C\right|+1+\left|\left(\cup_{A\in\mathcal{A}_{2}}A\right)\setminus C\right|+1-1$
$<\left|\mathcal{A}_{1}\right|+\left|\mathcal{A}_{2}\right|-1=\left|\left(\mathcal{A}_{1}\cup\mathcal{A}_{2}\right)\setminus\left\{C\right\}\right|$
@@
We notice that in all the observed cases, as suggested in [6], we have $\delta\mathcal{M}$ a maximal adjoint of $\mathcal{M}$ in the weak order of matroids. Using the tools developed in [11], we can construct a maximal matroid structure on a set of circuits if some conditions hold. We start by introducing the construction, let $\mathcal{X}\subseteq P(E)$ for the ground set $E$. A $\mathcal{X}$-matroid is a matroid $\mathcal{M}$ on the ground set $E$ such that $\mathcal{X}\subseteq C(\mathcal{M})$. [11] introduced the following upper bound on the rank function on a $\mathcal{X}$matroid.
@@ 6.7. Lemma:
Let $\mathcal{M}$ be an $\mathcal{X}$-matroid and $F\subseteq E$. Then $rank_{\mathcal{M}}(F)\leq val(F,S)$ for any proper $\mathcal{X}$-sequence $\mathcal{A}$. Furthermore, if the equality holds, then $rank_{\mathcal{M}}(F\setminus\left\{e\right\})=rank_{\mathcal{M}}(F)-1$ for all $e\in F\setminus\cup_{X\in\mathcal{A}}X$ and $rank_{\mathcal{M}}(F\cup\left\{e\right\})=rank_{\mathcal{M}}(F)$ for all $e\in\cup_{X\in\mathcal{A}}X$.
@@ 6.8. Theorem:
[…]
In our case we saw in theorem [5.5] that every collection of circuits of the form $\mathcal{C}_{\mathcal{C}_{B},C}$ is a circuit for every adjoint matroid, therefore the natural choice is $\mathcal{X}=\cup_{B\in\mathcal{M}}\left\{\mathcal{C}_{\mathcal{C}_{\mathcal{B}},C}\mid C\notin\mathcal{C}_{\mathcal{B}}\right\}$ making every adjoint of $\mathcal{M}$ an $\mathcal{X}$-matroid. Moreover, if $B\in\mathcal{M}$ and $\mathcal{A}=\left\{C\notin\mathcal{C}_{B}\right\}$, we have:
$val_{\mathcal{X}}\left(C(\mathcal{M})\right)\leq val(C(\mathcal{M}),\mathcal{A})=\left|C(\mathcal{M})\right|-\left|C(\mathcal{M})\setminus\mathcal{C}_{B}\right|=\left|\mathcal{C}_{B}\right|=rank(\mathcal{M}^{*})$
Therefore $val_{\mathcal{X}}\left(\mathcal{A}\right)\leq rank(\mathcal{M}^{*})$ for every $\mathcal{A}\subseteq C(\mathcal{M})$.
@@ 6.9. Lemma:
As $val_{\mathcal{X}}(\mathcal{A})\leq val_{\mathcal{X}}(\mathcal{A}\cup\left\{C\right\})\leq val_{\mathcal{X}}(\mathcal{A})+1$ for every $\mathcal{A}\in 2^{C(\mathcal{M})}$ and $C\in C(\mathcal{M})$ it is enough to prove the inequality for $\mathcal{D}\in(2^{C(\mathcal{M})}\setminus\mathcal{S}(\mathcal{M}))$. If $rank_{\mathcal{M}^{*}}(\cap_{C\in\mathcal{D}}C^{c})=0$, we must have $\left|\mathcal{D}\right|>rank(\mathcal{M}^{*})$ and the lemma follows from the above discussion. Otherwise, there exists an independent set $I$ of $\mathcal{M}^{*}$ of size $rank_{\mathcal{M}^{*}}(\cap_{C\in\mathcal{D}}C^{c})$ such that $I\cap C=\emptyset$ for all $C\in\mathcal{D}$. Let $B$ be a basis of $\mathcal{M}$ contained in the complement of $I$ and $\mathcal{A}=\left\{\mathcal{C}_{\mathcal{C}_{B},C}\mid C\in\mathcal{D}\right\}$. As $\mathcal{A}$ is a proper $\mathcal{X}$-sequence (in any order), we have the following:
@@
|$\left|\left\{C_{B,e}\mid e\in C\setminus B,\,C\in\mathcal{D}\right\}\right|=\left|B^{c}\right|-\left|I\right|=rank_{\mathcal{M}^{*}}(E)-rank_{\mathcal{M}^{*}}(\cap_{C\in\mathcal{D}}C^{c})$
@@ 6.10. Corollary:
Finally, we show that, with a small change in its construction, we have that $\delta\mathcal{M}$ is not smaller, in the weak order of matroids, then any adjoint of $\mathcal{m}$. Consequently, if it is an adjoint, then it is a maximal adjoint. Furthermore, using theorem [6.8], we see that if $val_{\mathcal{X}}$ is a rank function of a matroid, then it is the rank function of the small change of $\delta\mathcal{M}$.
@@ 6.11. Definition:
The small change to the derived matroid $\delta^{\prime}\mathcal{M}$ is defined to be the matroid on the set of circuits $C(\mathcal{M})$ and dependent sets $D^{(rank_{\mathcal{M}^{*}}(E))}(2^{C(\mathcal{M})}\setminus\mathcal{S})$, with $D{(k+1)}(\mathcal{A})=D(\mathcal{A}\cup(D{(k)}(\mathcal{A}))_{\leq k})$ and $D^{(0)}(\mathcal{A})=\mathcal{A}$.
We can see that the only change to $\delta\mathcal{M}$ is that in $\delta\mathcal{M}$ two dependent sets can have an independent intersection, which will become dependent in a later iteration. Let $\mathcal{D}$ be a minimal (in size and in the minimal number of iterations required to obtain it) dependent set in $\delta^{\prime}\mathcal{M}$ that is independent of an adjoint $\mathcal{M}^{\triangle}$. It must be the case that $\mathcal{D}=(\mathcal{D}_{1}\cup\mathcal{D}_{2})\setminus\left\{C\right\}$ with $\mathcal{D}_{1},\mathcal{D}_{2}$ dependent in $\mathcal{M}^{\triangle}$. Therefore, $\mathcal{D}_{1}\cap\mathcal{D}_{2}$ is dependent in $\mathcal{M}^{\triangle}$ and $\delta^{\prime}\mathcal{M}\nleq\mathcal{M}^{\triangle}$, which results in $\delta^{\prime}\mathcal{M}$ not being smaller than any adjoint of $\mathcal{M}$. We finish with a conjecture that will follow from Conjecture 1.3 in [11].
@@ Appendix A Computations
Using Theorem [5.6] we can construct a simple algorithm to check whether a matroid in the set of circuits of another matroid is its adjoint. We then implement this algorithm in Python using the Sage matroid package to calculate the list of adjoints of some matroids with a small number of circuits.
@@
$M,N$ are matroids and groundset($N$)=circuits($M$)
$N$ is an adjoint of $M$
for B in bases(M) do $\triangleright$ Checking if every family of fundamental circuits of a base $B$ is a base of $N$
@@
for B in bases(M) do $\triangleright$ Checking if every fundamental circuits of $(\mathcal{L}(M^{*}))^{opp}$ is a circuit of $N$
@@
We start by looking at corank $3$ examples for which the maximal adjoint $\mathcal{S}(\mathcal{M})$ is always an adjoint.
@@ A.2. Fano dual
  • $2$ isomorphic adjoints with $137$ bases.
@@
  • $6$ isomorphic adjoints with $136$ bases.
@@
  • $20$ adjoints with $138$ bases of two isomorphism types, $12$ of the first type and $8$ of the seconed type.
@@
  • $6$ isomorphic adjoints with $140$ bases.
@@
  • $6$ isomorphic adjoints with $239$ bases.
@@
  • $20$ adjoints with $241$ bases of two isomorphism types, $12$ of the first type and $8$ of the seconed type.
@@
  • $6$ isomorphic adjoints with $243$ bases.

### v3

@@ 6.4. Theorem
$1=\rho_{\mathcal{M}^{*}}(C_{1}^{c}\cap C_{2}^{c}\cap C_{3}^{c})=\rho_{\mathcal{M}^{*}}(C_{1}^{c}\cap C_{2}^{c})=\rho_{\mathcal{M}^{*}}(C_{1}^{c}\cap C_{3}^{c})=\rho_{\mathcal{M}^{*}}(C_{2}^{c}\cap C_{3}^{c})$
A set of size three in $\epsilon(2^{C(\mathcal{M})}\setminus\mathcal{S}(\mathcal{M}))$ is of the form $\left\{C_{2},C_{3},C_{4}\right\}$ with some $C_{1}\in\mathcal{C}(\mathcal{M})$ such that $\left\{C_{1},C_{2},C_{3}\right\},\left\{C_{1},C_{2},C_{4}\right\}\in 2^{C(\mathcal{M})}\setminus\mathcal{S}(\mathcal{M})$, which is again in $2^{C(\mathcal{M})}\setminus\mathcal{S}(\mathcal{M})$. As we also have $(\uparrow(2^{C(\mathcal{M})}\setminus\mathcal{S}(\mathcal{M})))_{3}=(2^{C(\mathcal{M})}\setminus\mathcal{S}(\mathcal{M}))_{3}$ and the maximal $\rho_{\mathcal{M}^{*}}$ of a set is $3$ we get $D(\delta\mathcal{M})=2^{C(\mathcal{M})}\setminus\mathcal{S}(\mathcal{M})$,
@@ 6.5. Theorem
$n-k-\rho_{\mathcal{M}^{*}}(\cap_{A\in\mathcal{A}_{i}}A)=\left|\left(\cup_{A\in\mathcal{A}_{i}}A\right)\setminus C\right|+1<\left|\mathcal{A}_{i}\right|$
@@
$n-k-\rho_{\mathcal{M}^{*}}(\cap_{A\in\left(\mathcal{A}_{1}\cup\mathcal{A}_{2}\right)\setminus\left\{C\right\}}A)\leq n-k-\rho_{\mathcal{M}^{*}}(\cap_{A\in\mathcal{A}_{1}\cup\mathcal{A}_{2}}A)$
$\left|\left(\cup_{A\in\mathcal{A}_{1}\cup\mathcal{A}_{2}}A\right)\setminus C\right|+1\leq\left|\left(\cup_{A\in\mathcal{A}_{1}}A\right)\setminus C\right|+1+\left|\left(\cup_{A\in\mathcal{A}_{2}}A\right)\setminus C\right|+1-1$
$\left|\mathcal{A}_{1}\right|+\left|\mathcal{A}_{2}\right|-1=\left|\left(\mathcal{A}_{1}\cup\mathcal{A}_{2}\right)\setminus\left\{C\right\}\right|$
@@
We notice that in all observed cases, as suggested in [6], we have $\delta\mathcal{M}$ a maximal adjoint of $\mathcal{M}$ in the weak order of matroids. Using the tools developed in [11], we can construct a maximal matroid structure on a set of circuits if some conditions hold. We start by introducing the construction, let $\mathcal{X}\subseteq P(E)$ for the ground set $E$. A $\mathcal{X}$-matroid is a matroid $\mathcal{M}$ on the ground set $E$ such that $\mathcal{X}\subseteq C(\mathcal{M})$. [11] introduced the following upper bound on the rank function on a $\mathcal{X}$ -matroid.
@@ 6.7. Lemma:
Let $\mathcal{M}$ be an $\mathcal{X}$-matroid and $F\subseteq E$. Then $\rho_{\mathcal{M}}(F)\leq val(F,S)$ for any proper $\mathcal{X}$-sequence $\mathcal{A}$. Furthermore, if the equality holds, then $\rho_{\mathcal{M}}(F\setminus\left\{e\right\})=\rho_{\mathcal{M}}(F)-1$ for all $e\in F\setminus\cup_{X\in\mathcal{A}}X$ and $\rho_{\mathcal{M}}(F\cup\left\{e\right\})=\rho_{\mathcal{M}}(F)$ for all $e\in\cup_{X\in\mathcal{A}}X$.
@@ 6.8. Theorem:
$val_{\mathcal{X}}\left(F\right)=min\left\{val\left(F,\mathcal{A}\right)\mid\mathcal{A}\text{ is a proper }\mathcal{X}\text{-sequence}\right\}$
In our case, we saw in Theorem 5.7 that every collection of circuits of the form $\mathcal{C}_{\mathcal{C}_{B},C}$ is a circuit for every adjoint matroid, therefore the natural choice is $\mathcal{X}=\cup_{B\in\mathcal{M}}\left\{\mathcal{C}_{\mathcal{C}_{\mathcal{B}},C}\mid C\notin\mathcal{C}_{\mathcal{B}}\right\}$ making every adjoint of $\mathcal{M}$ an $\mathcal{X}$-matroid. Moreover, if $B\in\mathcal{M}$ and $\mathcal{A}=\left\{C\notin\mathcal{C}_{B}\right\}$, we have:
$val_{\mathcal{X}}\left(C(\mathcal{M})\right)\leq val(C(\mathcal{M}),\mathcal{A})=\left|C(\mathcal{M})\right|-\left|C(\mathcal{M})\setminus\mathcal{C}_{B}\right|=\left|\mathcal{C}_{B}\right|=\rho(\mathcal{M}^{*})$
Therefore $val_{\mathcal{X}}\left(\mathcal{A}\right)\leq\rho(\mathcal{M}^{*})$ for every $\mathcal{A}\subseteq C(\mathcal{M})$.
@@ 6.9. Lemma:
As $val_{\mathcal{X}}(\mathcal{A})\leq val_{\mathcal{X}}(\mathcal{A}\cup\left\{C\right\})\leq val_{\mathcal{X}}(\mathcal{A})+1$ for every $\mathcal{A}\in 2^{C(\mathcal{M})}$ and $C\in C(\mathcal{M})$ it is enough to prove the inequality for $\mathcal{D}\in(2^{C(\mathcal{M})}\setminus\mathcal{S}(\mathcal{M}))$. If $\rho_{\mathcal{M}^{*}}(\cap_{C\in\mathcal{D}}C^{c})=0$, we must have $\left|\mathcal{D}\right|>\rho(\mathcal{M}^{*})$ and the lemma follows from the above discussion. Otherwise, there exists an independent set $I$ of $\mathcal{M}^{*}$ of size $\rho_{\mathcal{M}^{*}}(\cap_{C\in\mathcal{D}}C^{c})$ such that $I\cap C=\emptyset$ for all $C\in\mathcal{D}$. Let $B$ be a basis of $\mathcal{M}$ contained in the complement of $I$ and $\mathcal{A}=\left\{\mathcal{C}_{\mathcal{C}_{B},C}\mid C\in\mathcal{D}\right\}$. As $\mathcal{A}$ is a proper $\mathcal{X}$-sequence (in any order), we have the following:
@@
$\left|\left\{C_{B,e}\mid e\in C\setminus B,\,C\in\mathcal{D}\right\}\right|=\left|B^{c}\right|-\left|I\right|=\rho_{\mathcal{M}^{*}}(E)-\rho_{\mathcal{M}^{*}}(\cap_{C\in\mathcal{D}}C^{c})$
@@ 6.10. Corollary:
Finally, we demonstrate that, with a slight modification in its construction, $\delta\mathcal{M}$ is not smaller than any adjoint of $\mathcal{M}$ in the weak order of matroids. Consequently, if $\delta\mathcal{M}$qualifies as an adjoint, it must be the maximal adjoint. Moreover, applying Theorem 6.8, we establish that if $val_{\mathcal{X}}$ s a rank function of a matroid, then it serves as the rank function of the modified version of $\delta\mathcal{M}$.
@@ 6.11. Definition:
The small change to the combinatorial derived matroid $\delta^{\prime}\mathcal{M}$ is defined to be the matroid on the set of circuits $C(\mathcal{M})$ and dependent sets $D^{(\rho_{\mathcal{M}^{*}}(E))}(2^{C(\mathcal{M})}\setminus\mathcal{S})$, with $D{(k+1)}(\mathcal{A})=D(\mathcal{A}\cup(D{(k)}(\mathcal{A}))_{\leq k})$ and $D^{(0)}(\mathcal{A})=\mathcal{A}$.
We can see that the only change to $\delta\mathcal{M}$ is that in $\delta\mathcal{M}$ two dependent sets can have an independent intersection, which eventually becomes dependent in a later iteration. Let $\mathcal{D}$ be a minimal, both in size and in the number of iterations required to obtain it, that remains independent in an adjoint $\mathcal{M}^{\triangle}$. It must be the case that $\mathcal{D}=(\mathcal{D}_{1}\cup\mathcal{D}_{2})\setminus\left\{C\right\}$ with $\mathcal{D}_{1},\mathcal{D}_{2}$ dependent in $\mathcal{M}^{\triangle}$. Therefore, $\mathcal{D}_{1}\cap\mathcal{D}_{2}$ is dependent in $\mathcal{M}^{\triangle}$ and $\delta^{\prime}\mathcal{M}\nleq\mathcal{M}^{\triangle}$, which results in $\delta^{\prime}\mathcal{M}$ not being smaller than any adjoint of $\mathcal{M}$. We finish with a conjecture that will follow from Conjecture 1.3 in [11].
@@ Appendix A Computations
Using Theorem 5.2 we can construct a simple algorithm to check whether a matroid in the set of circuits of another matroid is its adjoint. We then implement this algorithm in Python using the Sage matroid package to calculate the list of adjoints of some matroids with a small number of circuits.
@@
$M,N$ are matroids and groundset(NN)=circuits($M$)
NN is an adjoint of $M$
for B in bases(M) do $\triangleright$ Checking if every family of fundamental circuits of a base $B$ is a base of NN
@@
for B in bases(M) do $\triangleright$ Checking if every fundamental circuits of $(\mathcal{L}(M^{*}))^{opp}$ is a circuit of NN
@@
We start by looking at co-rank $3$ examples for which the maximal adjoint $\mathcal{S}(\mathcal{M})$ is always an adjoint.
@@ A.2. Fano dual
  • 22 isomorphic adjoints with $137$ bases.
@@
  • 66 isomorphic adjoints with $136$ bases.
@@
  • 2020 adjoints with $138$ bases of two isomorphism types, $12$ of the first type and $8$ of the seconed type.
@@
  • 66 isomorphic adjoints with $140$ bases.
@@
  • 66 isomorphic adjoints with $239$ bases.
@@
  • 2020 adjoints with $241$ bases of two isomorphism types, $12$ of the first type and $8$ of the seconed type.
@@
  • 66 isomorphic adjoints with $243$ bases.